OpenStudy (anonymous):

A car with a mass of 800kg experience a friction force of 180N and travels at a constant speed of 54km/h. Calculate the power required from the engine to travel at the constant speed.

OpenStudy (anonymous):

F * v

OpenStudy (anonymous):

dont you need distance and time to calculate power?

OpenStudy (anonymous):

yes you do :) velocity = distance/time. so both are involved :)

OpenStudy (anonymous):

no need distance...... its enough... bcoz 180N is given... thn what else needed

OpenStudy (anonymous):

power=work done/time

OpenStudy (anonymous):

work done=force x distance

OpenStudy (anonymous):

p=w/t w=fd p=fd/t d/t= v so p=fv you have both